### v1

# A stochastic method to estimate a zero-inflated two-part mixed model for human microbiome data

Abstract
Human microbiome studies based on genetic sequencing techniques produce compositional longitudinal data of the relative abundances of microbial taxa over time, allowing to understand, through mixed-effects modeling, how microbial communities evolve in response to clinical interventions, environmental changes, or disease progression. In particular, the Zero-Inflated Beta Regression (ZIBR) models jointly and over time the presence and abundance of each microbe taxon, considering the compositional nature of the data, its skewness, and the over-abundance of zeros. However, as for other complex random effects models, maximum likelihood estimation suffers from the intractability of likelihood integrals. Available estimation methods rely on log-likelihood approximation, which is prone to potential limitations such as biased estimates or unstable convergence. In this work we develop an alternative maximum likelihood estimation approach for the ZIBR model, based on the Stochastic Approximation Expectation Maximization (SAEM) algorithm. The proposed methodology allows to model unbalanced data, which is not always possible in existing approaches. We also provide estimations of the standard errors and the log-likelihood of the fitted model. The performance of the algorithm is established through simulation, and its use is demonstrated on two microbiome studies, showing its ability to detect changes in both presence and abundance of bacterial taxa over time and in response to treatment.
Keywords compositional data; longitudinal data; microbiome data; SAEM algorithm; ZIBR model.
1 Introduction
The human microbiome, a complex community of microorganisms, plays a crucial role in the body’s functions and health. It influences metabolic pathways, drug metabolism, and contributes to the bioconversion of nutrients, detoxification, and protection against pathogens (Dekaboruah et al.,, 2020). In particular, the gut microbiota has been shown to interact with the host immune system, influencing the development of some diseases (Clemente et al.,, 2012). The microbiome significantly impacts the human genome, with genotypes influencing its composition and activity, while the microbiome alters the expression of genetic risk for chronic inflammatory and immune conditions (Jeyakumar et al.,, 2019).
Given its major role in human health, it is crucial to have an accurate and reliable framework for the collection and analysis of human microbiome data. Current collection methods include high-throughput sequencing technologies such as the 16S ribosomal RNA (rRNA) sequencing approach and shotgun sequencing. These benchmark procedures are at the basis of leading studies such as the Human Microbiome Project (Turnbaugh et al.,, 2007), which characterized microbiomes in five major parts of the human body: airway, skin, oral cavity, digestive tract, and vagina. They produce sequence counts that are not comparable between samples, so traditional approaches consist in normalizing by the total sequence count, producing compositional data (Tyler et al.,, 2014). If multiple observations are recorded at different points in time, an additional component is introduced, leading to the analysis of longitudinal compositional data.
Regarding data analysis, Kodikara et al., (2022) compiled some recent models developed to study longitudinal microbiome data from sample sequencing, considering those that model count data and those that model relative abundances. One of the latter is the Zero-Inflated Beta Regression (ZIBR) (Chen and Li,, 2016), which is defined as a two-stage mixed effects model which is based on the work of Ospina and Ferrari, (2012), that allows the inclusion of clinical covariates, both to explain the presence or absence of a certain bacterial taxon and, in case of presence, the influence of these covariates on the relative abundance of the taxon. It also provides a comprehensive approach to analyse longitudinal compositional microbiome data taking into account its bounded nature, skewness and the over-abundance of zeros. Since its appearance, ZIBR has been succesfully applied in several studies (Hu et al.,, 2022; D’Agata et al.,, 2019) as it is capable of treating the features above mentioned, explaining within-subject correlations and providing methods to conduct hypothesis tests on the significance of covariates.
As for other complex mixed-effects models, the Maximum Likelihood (ML) estimation method for ZIBR proposed by its authors relies on approximating the log-likelihood using Gauss-Hermite quadrature and numerical optimization of this expression. However, there is evidence that in certain scenarios the estimators obtained in this way may be biased and, in the case of generalized mixed models, even be outperformed by other techniques (Handayani et al.,, 2017). Additionally, the proposed estimation method can only be used for balanced data; that is, with the same number of observations per individual. In clinical studies this is not often the case, resulting in analysis with potentially misleading conclusions (Powney et al.,, 2014). Therefore, new strategies are required to address the challenges posed by missing data (Myers,, 2000). The ZIBR model can also be thought of a particular case of the Generalized Additive Model for Location, Scale and Shape (GAMLSS, Rigby and Stasinopoulos,, 2005) and its parameters can be estimated in this framework. Although in this case the available maximum estimation strategy does allow to handle missing data, it still relies on a penalized local log-likelihood approximation which can cause problems in performing likelihood ratio tests, particularly in complex GLMMs such as the ZIBR (see Stasinopoulos et al., (2017)).
All things considered, the versatile features of the ZIBR model make it a promising choice for the analysis of compositional microbiome data, despite possible drawbacks in existing estimation strategies. Therefore, for the final purpose of precisely identifying those taxa responsive to disease onset, changes in environmental conditions or specific interventions, any possible improvement in the estimation method which is able to provide more accurate estimates will amount to significant progress in the understanding of human microbiome and its relation to human health.
Following this aim, in this work we propose a new estimation framework for the ZIBR model on longitudinal compositional data based on the Stochastic Approximation EM (SAEM) algorithm (Delyon et al.,, 1999). This algorithm provides an exact maximum likelihood estimation strategy in missing data models for which the EM algorithm (Dempster et al.,, 1977) is not directly applicable because the complexity of the likelihood function does not allow for exact calculation of its conditional expectation. This would be the case of the ZIBR model. The SAEM algorithm not only preserves the good behaviour of the EM algorithm in terms of convergence, unbiasedness and monotonicity, but has also shown interesting properties in complex mixed models (Márquez et al.,, 2023; Meza et al.,, 2012) and includes procedures for statistical inference and hypothesis testing (Samson et al.,, 2007). Furthermore, it can be combined with MCMC techniques for improved performance (Kuhn and Lavielle,, 2005) and can be extended to Restricted Maximum Likelihood (REML) estimation (Meza et al.,, 2007).
In this article we extend the algorithm to distributions not belonging to the exponential family and derive the explicit expressions at all its steps, for both parameter estimation and log-likelihood approximation, once the ML estimators have been obtained. We also obtain approximations of the standard errors of the estimators, by means of the stochastic approximation of the Fisher information matrix. This allows us to provide a comprehensive estimation approach that avoids downsides related to likelihood approximations, is able to incorporate unbalanced data, and facilitates the inference pipeline, from modeling to covariate effects testing, under the same framework.
The structure of the document is as follows: in Section 2, we introduce the ZIBR model and develop the SAEM based inference method to be used in our work. In Section 3 we present simulation studies on synthetic data generated under different settings, comparing the results obtained with our approach and those given by estimation based on likelihood approximation or penalization, and in Section 4 we assess the behaviour of the proposed routine on a dataset coming from clinical microbiome studies. Finally, Section 5 closes the article with the main conclusions, a discussion of the results, and possible limitations and future developments.
2 Models and methods
In this section we describe the ZIBR model for longitudinal compositional data (Chen and Li,, 2016), we revise the foundations of the SAEM algorithm for parameter estimation by maximum likelihood as well as log-likelihood estimation through Importance Sampling, and present its extension to the ZIBR model.
2.1 The ZIBR model for longitudinal compositional data
The ZIBR model describes the presence and abundance of a single bacterial taxon on different individuals over time, and can be subsequently applied to different bacteria. Let $y_{it}$ be the relative abundance of a bacterial taxon in the individual $i$ at time $t$, $1\leq i\leq N$, $1\leq t\leq T_{i}$. The model assumes that $y_{it}$ follows the distribution:
$y_{it}\sim\begin{cases}0&\mbox{with prob. }1-p_{it},\\ Beta(u_{it}\phi,(1-u_{it})\phi)&\mbox{with prob. }p_{it}\\ \end{cases}$ (1)
with $\phi>0$ and $10<u_{it},p_{it}<$ 1. These two last components are characterized by
[…]
where $a_{i}$ and $b_{i}$ are individual specific intercepts, $\alpha$ and $\beta$ are vectors of regression coefficients and $X_{it}$ and $Z_{it}$ are covariates for each individual and time point. We further consider that each one of the random intercepts follows a normal distribution, independently from each other:
$a_{i}\sim N(a,\sigma^{2}_{1}),\hskip 14.22636ptb_{i}\sim N(b,\sigma^{2}_{2}).$
From Equations 1 and 2, it can be seen that the ZIBR model explicitly includes a component that is responsible for the presence of zeros in the data. It is also clear that conveniently defined covariates $X_{it}$ and $Z_{it}$ can influence both the probability of presence or absence of a bacterial taxon (through the logistic regression that defines $p_{it}$) and the magnitude of its relative abundance (through the $u_{it}$ component in the proposed Beta distribution). Furthermore, the inclusion of a random intercept allows modeling correlations in observations from the same individual. Even though it is easy to expand the definition to consider random slopes, in practice it is enough to consider just a random intercept (Min and Agresti,, 2005).
The model parameter $\theta=(\phi,a,b,\alpha,\beta,\sigma^{2}_{1},\sigma^{2}_{2})$ can be estimated by maximum likelihood. From Equations 1 and 2, the likelihood function for data $\bm{y}=\left(y_{it},1\leq i\leq N,1\leq t\leq T_{i}\right)$ is
$L(\theta;\bm{y})=\prod_{i=1}^{N}\int_{\mathbb{R}}\int_{\mathbb{R}}\prod_{t=1}^{T_{i}}(1-p_{it})^{\mathbbm{1}_{\left\{y_{it}=0\right\}}}\left[p_{it}f(y_{it};u_{it},\phi)\right]^{\mathbbm{1}_{\left\{y_{it}>0\right\}}}g\left(a_{i},b_{i}|a,\sigma^{2}_{1},b,\sigma^{2}_{2}\right)da_{i}db_{i}$ (3)
where $f(y_{it};u_{it},\phi)$ is the Beta density function with parameters $u_{it}$ and $\phi$ on $y_{it}$:
$f(y_{it};u_{it},\phi)=\frac{\Gamma(\phi)}{\Gamma(u_{it}\phi)\Gamma((1-u_{it})\phi)}y_{it}^{u_{it}\phi-1}(1-y_{it})^{(1-u_{it})\phi-1},$
and $g$ is the product of the two univariate normal density functions of random effects $a_{i}$ and $b_{i}$.
Given the impossibility of analytical calculation of the integral shown in Equation 3, an approximation can be achieved by means of the Gauss-Hermite quadrature (GHQ). With this approximation, and through numerical optimization, the maximum likelihood estimators for $\theta$ can be found as proposed by Chen and Li, (2016). Hypothesis tests for the significance of covariates can also be conducted, in particular the Likelihood Ratio Test (LRT). The implementation of this approach is available in the ZIBR package (Zhang Chen,, 2023) developed for the R software. In addition to this alternative, the gamlss package (Rigby and Stasinopoulos,, 2005) can also be used, which, in a similar manner to the aforementioned one, is based on a penalized quasilikelihood approximation and its optimization based on numerical algorithms (Rigby and Stasinopoulos,, 1996). A notable advantage of this implementation, however, is its capacity to handle unbalanced data, which renders it a suitable option for a comparative analysis of the results obtained in this study.
2.2 The SAEM algorithm for mixed effects models
The Stochastic Approximation Expectation-Maximization (SAEM) algorithm (Delyon et al.,, 1999) is a powerful tool for estimating population parameters in complex mixed effect models. This algorithm is applicable for the iterative computation of ML estimates in a wide variety of incomplete data statistical problems in which the Expectation step of the EM algorithm is not explicit; in particular in mixed effects models, where the individual random effects are treated as non-observed data. Let $\bm{y}=\left(y_{it},1\leq i\leq N,1\leq t\leq T_{i}\right)$ and $\bm{\varphi}=\left(\varphi_{i},1\leq i\leq N\right)$ denote the observed and non-observed data, respectively, so the complete data of the model are $(\bm{y},\bm{\varphi})$. In this case, the SAEM algorithm consists of replacing the usual E-step of EM with a stochastic approximation procedure. Given an initial point $\theta^{(0)}$, iteration $q$ of the algorithm writes:
  • Simulation (S) step: Draw a realization $\bm{\varphi}^{(q)}$ from the conditional distribution $p\left(\cdot|\;\bm{y};\theta^{(q-1)}\right)$.
  • Stochastic Approximation (SA) step: Update $s_{q}(\theta)$, the approximation of the conditional expectation $\mathbb{E}\left[\log p\left(\bm{y},\bm{\varphi}^{(q)};\theta\right)|\bm{y},\theta^{(q-1)}\right]$:
    $s_{q}(\theta)=s_{q-1}(\theta)+\gamma_{q}\left(\log p\left(\bm{y},\bm{\varphi}^{(q)};\theta\right)-s_{q-1}(\theta)\right)$
    where $\{\gamma_{q}\}_{q\in\mathbb{N}}$ is a decreasing sequence of stepsizes with $\gamma_{1}=1$.
  • Maximization (M) step: Update $\theta^{(q)}$ according to $\theta^{(q)}=\operatorname*{arg\,max}_{\theta}s_{q}(\theta).$
There are some important remarks on the the working details of the SAEM algorithm. In the case of complex mixed effects models, such as ZIBR, the conditional distribution of the non-observed data $p\left(\cdot|\;\bm{y};\theta^{(q-1)}\right)$ cannot be computed in closed form and simulation from it cannot be carried out directly (Kuhn and Lavielle,, 2005). However, a MCMC approach can be used in the Simulation step of the SAEM algorithm described above, consisting in applying the Metropolis-Hastings algorithm (Metropolis et al.,, 1953) with different proposal kernels, in order to approximate $p\left(\cdot|\;\bm{y};\theta^{(q-1)}\right)$ with a Markov chain with defined transition probabilities.
Also, convergence can be improved by generating more than one Markov chain or realization at simulation and by applying a Monte Carlo scheme. That is, at the Simulation step $m$ realizations $\bm{\varphi}^{(q,l)}\sim p\left(\cdot|\;\bm{y};\theta^{(q-1)}\right)$, $1\leq l\leq m$, are drawn, and in the SA step the approximation of the conditional expectation is updated as
$s_{q}(\theta)=s_{q-1}(\theta)+\gamma_{q}\left(\frac{1}{m}\sum_{l=1}^{m}\log p\left(\bm{y},\bm{\varphi}^{(q,l)};\theta\right)-s_{q-1}(\theta)\right).$
If the complete-data model belongs to the exponential family, that is, if
$\log p(\bm{y},\bm{\varphi};\theta)=-\Psi(\theta)+\langle S(\bm{y},\bm{\varphi}),\xi(\theta)\rangle$
where $S(\bm{y},\bm{\varphi})$ represents a sufficient statistic of the data, then, the SA step reduces to:
$F_{q}=F_{q-1}+\gamma_{q}\left(\frac{1}{m}\sum_{l=1}^{m}S(\bm{y},\bm{\varphi}^{(q,l)})-F_{q-1}\right)$ (4)
and $s_{q}(\theta)=-\Psi(\theta)+\langle F_{q},\xi(\theta)\rangle$; that is, the actualization is made only on the sufficient statistic. This scheme can be applied even to models outside the exponential family, provided that a part of the model belongs to this family. However, we cannot speak of updating a sufficient statistic, but rather of a data summary function. Under general circumstances (Delyon et al.,, 1999; Kuhn and Lavielle,, 2005), the convergence of the parameter sequence $\theta^{(q)}$ toward a (local) maximum of the likelihood $\hat{\theta}$ is guaranteed, regardless of the starting point $\theta^{(0)}$ (Celeux et al.,, 1995).
The sequence of stepsizes $\{\gamma_{q}\}_{q\in\mathbb{N}}$ is usually set to 1 during the first iterations to avoid getting stuck in local maxima. In this way the first iterations are identical to those of the Monte Carlo EM (MCEM) algorithm (Wei and Tanner,, 1990), which is known for its slow convergence rate. To avoid this scenario, in later iterations of SAEM $\gamma_{q}$ decreases to zero to force convergence with fewer iterations. Details of application of the SAEM algorithm to complex mixed-effects models can be found in Meza et al., (2012); Márquez et al., (2023); Arribas-Gil et al., (2014) and de la Cruz et al., (2024).
2.3 The SAEM algorithm for ZIBR parameter estimation
[…]
$\begin{split}p(\bm{y},\bm{\varphi};\theta)=&p(\bm{y}|\bm{\varphi};\alpha,\beta,\phi)p(\bm{\varphi}|\bm{\mu},\mathbf{G})\\ \propto&|\mathbf{G}|^{-\frac{N}{2}}\prod_{i}\exp{\left(-\frac{(\varphi_{i}-\bm{\mu})^{T}\mathbf{G^{-1}}(\varphi_{i}-\bm{\mu})}{2}\right)}\\ &\times\prod_{i,t}(1-p_{it})^{\mathbbm{1}_{\left\{y_{it}=0\right\}}}p_{it}^{\mathbbm{1}_{\left\{y_{it}>0\right\}}}f(y_{it};u_{it},\phi)^{\mathbbm{1}_{\left\{y_{it}>0\right\}}}.\end{split}$ (5)
Like most zero-inflated models, the ZIBR model cannot be considered part of the exponential family (Eggers,, 2015). However, the decomposition presented in Equation 5 allows us to propose a simplified structure for the SAEM algorithm (Equation 4). For the multivariate normal part corresponding to the random effects, the actualization in the SA step is done on the respective sufficient statistics. For the mixture distribution corresponding to the observed data, $\bm{y}|\bm{\varphi};\alpha,\beta,\phi$, maximization of the conditional log-likelihood is followed by estimates updates, as suggested for non-exponential family models (Comets et al.,, 2021). Then, the Maximum Likelihood iterative estimation algorithm for the parameters of the ZIBR model writes, for a given starting point $\theta^{(0)}$ and at iteration $q$, as:
  1. 1.
    Simulation step: draw $\varphi^{(q)}_{i},i=1,\cdots,N$ from the distribution $p(\cdot|\bm{y};\theta^{(q-1)})$.
  2. 2.
    Stochastic Approximation step: update the summary data functions $F_{1}(\bm{y},\bm{\varphi})$ and $F_{2}(\bm{y},\bm{\varphi})$ with the scheme:
    $\begin{split}F_{1}^{(q)}(\bm{y},\bm{\varphi})&=F_{1}^{(q-1)}(\bm{y},\bm{\varphi})+\gamma_{q}\left(\sum_{i}{\varphi_{i}^{(q)}}-F_{1}^{(q-1)}(\bm{y},\bm{\varphi})\right)\\ F_{2}^{(q)}(\bm{y},\bm{\varphi})&=F_{2}^{(q-1)}(\bm{y},\bm{\varphi})+\gamma_{q}\left(\sum_{i}{\varphi_{i}^{(q)}{\varphi_{i}^{(q)}}^{T}}-F_{2}^{(q-1)}(\bm{y},\bm{\varphi})\right).\end{split}$ (6)
    where $\{\gamma_{q}\}_{q\in\mathbb{N}}$ is a decreasing sequence of stepsizes with $\gamma_{1}=1$.
  3. 3.
    Maximization step: update the parameters of the model with
    $\begin{split}\bm{\mu}^{(q)}&=\frac{F_{1}^{(q)}(\bm{y},\bm{\varphi})}{N}\\ \mathbf{G}^{(q)}&=\frac{F_{2}^{(q)}(\bm{y},\bm{\varphi})}{N}-\frac{F_{1}^{(q)}(\bm{y},\bm{\varphi}){F_{1}^{(q)}(\bm{y},\bm{\varphi})}^{T}}{N^{2}}\end{split}$ (7)
    Given the form of the model definition in the Beta part, steps 2 and 3 are modified by first calculating
    $\left(\tilde{\beta}^{(q)},\tilde{\phi}^{(q)}\right)=$ (8)
    $\operatorname*{arg\,max}_{\beta,\phi}\sum_{i,t}$ $\left[\mathbbm{1}_{\left\{y_{it}>0\right\}}\left(\log{\frac{\Gamma(\phi)}{\Gamma\left(u_{it}^{(q)}\phi\right)\Gamma\left(\left(1-u_{it}{(q)}\right)\phi\right)}}+u_{it}^{(q)}\phi\log{y_{it}}+\phi\left(1-u_{it}^{(q)}\right)\log{(1-y_{it})}\right)\right]$
    and
    $\tilde{\alpha}^{(q)}=\operatorname*{arg\,max}_{\alpha}\sum_{i,t}\left[\mathbbm{1}_{\left\{y_{it}>0\right\}}\log{\left(p_{it}^{(q)}\right)}+\mathbbm{1}_{\left\{y_{it}=0\right\}}\log{\left(1-p_{it}^{(q)}\right)}\right].$ (9)
    where $u_{it}^{(q)}=u_{it}^{(q)}(b_{i},\beta)$ and $p_{it}^{(q)}=p_{it}^{(q)}(a_{i},\alpha)$ are calculated using $\varphi^{(q)}_{i}$ and Equation 2. Maximization in (8) and (9) is achieved numerically. Finally, the values are updated by doing
    $\begin{split}\phi^{(q)}&=\phi^{(q-1)}+\gamma_{q}\left(\tilde{\phi}^{(q)}-\phi^{(q-1)}\right)\\ \alpha^{(q)}&=\alpha^{(q-1)}+\gamma_{q}\left(\tilde{\alpha}^{(q)}-\alpha^{(q-1)}\right)\\ \beta^{(q)}&=\beta^{(q-1)}+\gamma_{q}\left(\tilde{\beta}^{(q)}-\beta^{(q-1)}\right)\end{split}$ (10)
Let us discuss the details of this implementation. As mentioned in 2.2, the choice of the starting point $\theta^{(0)}$ for SAEM does not affect its convergence; however, it is recommended to use values obtained in previous studies or with other estimation methods. Following the example of the existing implementation of the saemix packageComets et al., (2017), we will use $\gamma_{q}$ defined as follows:
$\gamma_{q}=\begin{cases}1&\mbox{if $q\leq K_{1}$,}\\ \frac{1}{q-K_{1}}&\mbox{if $K_{1}<q\leq K_{1}+K_{2}$.}\\ \end{cases}$
where $K_{1}+K_{2}$ is the total number of iterations.
We also discussed in section 2.2 that it is possible to improve the performance of the algorithm by taking multiple sequences or Markov chains in the Simulation step, and using Monte Carlo in Equations 7 and 10. Furthermore, during the SA step, we obtain sequences that allow to estimate $\mathbb{E}\left(\varphi_{i}|y_{i};\hat{\theta}\right)$ and $\mbox{Var}\left(\varphi_{i}|y_{i};\hat{\theta}\right)$ to be calculated, values necessary to approximate the log-likelihood through Importance Sampling, with which the Likelihood Ratio Test (LRT) can be computed, as presented in the following subsection.
2.3.1 Approximation of the log-likelihood using Importance Sampling
The log-likelihood of the observed data cannot be computed in closed form for complex mixed effects models, but its estimation is required to perform the LRT and to compute information criteria for a given model. One approximation method is given by the application of the Importance Sampling algorithm (Kloek and van Dijk,, 1978). Let $\mathcal{L}\mathcal{L}_{y}(\hat{\theta})$ be the log-likelihood of the model at the vector of population parameter estimates, that is $\mathcal{LL}_{y}(\hat{\theta})=\log p(\bm{y};\hat{\theta})$ where $p(\bm{y};\hat{\theta})=L(\hat{\theta};\bm{y})$ is the joint probability distribution function of the observed data given $\hat{\theta}$. Notice that $\mathcal{LL}_{y}(\hat{\theta})=\log p(\bm{y};\hat{\theta})=\sum_{i=1}^{N}\log p(y_{i};\hat{\theta})$ and, for some proposal distribution $\tilde{p}_{\varphi_{i}}$ absolutely continuous with respect to $p_{\varphi_{i}}$, we have
[…]
That is, $p(y_{i};\hat{\theta})$ can be expressed as an expectation which can be approximated by:
  1. 1.
    Obtain a random sample of size $K$ $\varphi_{i}^{(1)},\cdots,\varphi_{i}^{(K)}$ from the proposal distribution $\tilde{p}_{\varphi_{i}}$;
  2. 2.
    Compute the empirical mean $\hat{p}_{(i,K)}=\frac{1}{K}\sum_{k=1}^{K}p(y_{i}|\varphi_{i}^{(k)};\hat{\theta})\frac{p(\varphi_{i}^{(k)};\hat{\theta})}{\tilde{p}(\varphi_{i}^{(k)};\hat{\theta})}$
An optimal proposal distribution would be the conditional distribution $p_{\varphi_{i}|y_{i}}$ since in that case the estimator of the expectation has zero variance. But since the closed form expression of the distribution is not available, we choose a proposal close to this optimal distribution, based on the empirically estimated conditional mean and variance, $\mu_{i}=\hat{\mathbb{E}}\left[\varphi_{i}|y_{i};\hat{\theta}\right]$ and $\sigma^{2}_{i}=\hat{\mbox{Var}}\left[\varphi_{i}|y_{i};\hat{\theta}\right]$, of the simulated random effects during the simulation step of the SAEM algorithm. Then, the proposed candidate $\varphi_{i}^{(k)}$, with $k=1,\dots,K$, is drawn with a noncentral student $t$-distribution $\varphi_{i}^{(k)}=\mu_{i}+\sigma_{i}\times T_{i,k}$, with $T_{i,k}\sim t_{\nu}$ i.i.d., where $t_{\nu}$ denotes a Student’s t-distribution with $\nu$ degrees of freedom. In this work, unless otherwise mentioned, the parameters for calculating the log-likelihood will be $\nu=5$ and $K=500$.
2.3.2 Stochastic approximation of the standard errors
In addition to providing estimates of the parameters of a model, it is desirable that the estimation method is capable of also estimating its standard errors, with the objective of constructing confidence intervals or performing statistical tests on individual estimators, such as the Wald test. In the case of maximum likelihood estimation, these errors can asymptotically be calculated based on the Fisher information matrix of the model, which for complex models cannot be computed in a closed form. Based on the Louis’s missing information principle (Louis,, 1982) it is possible to compute an estimation of the Fisher information matrix. According to this principle, we have the identity:
$\partial_{\theta}^{2}\log p(\bm{y};\theta)=\mathbb{E}\left(\partial_{\theta}^{2}\log p(\bm{y},\bm{\varphi};\theta)|\bm{y};\theta\right)+\mbox{Cov}\left(\partial_{\theta}\log p(\bm{y},\bm{\varphi};\theta)|\bm{y};\theta\right)$
where
$\mbox{Cov}\left(\partial_{\theta}\log p(\bm{y},\bm{\varphi};\theta)|\bm{y};\theta\right)=$ $\mathbb{E}\left(\partial_{\theta}\log p(\bm{y},\bm{\varphi};\theta)\partial_{\theta}\log p(\bm{y},\bm{\varphi};\theta)^{T}|\bm{y};\theta\right)$
$-\mathbb{E}\left(\partial_{\theta}\log p(\bm{y},\bm{\varphi};\theta)|\bm{y};\theta\right)\mathbb{E}\left(\partial_{\theta}\log p(\bm{y},\bm{\varphi};\theta)|\bm{y};\theta\right)^{T}$
Given this, the second order derivative of the observed likelihood function with respect to parameter $\hat{\theta}$, $\partial_{\theta}^{2}L(\hat{\theta};\bm{y})$, can be approximated by the sequence $\{H_{q}\}_{q\in\mathbb{N}}$ which is calculated at iteration $q$ of the SAEM algorithm as:
$D_{q}$ $=$ $D_{q-1}+\gamma_{q}\left[\partial_{\theta}\log p(\bm{y},\bm{\varphi}^{(q)};\bm{\theta}^{(q)})-D_{q-1}\right]$
$G_{q}$ $=$ $G_{q-1}+\gamma_{q}\left[\partial^{2}_{\theta}\log p(\bm{y},\bm{\varphi}^{(q)};\bm{\theta}^{(q)})\right.$
$\left.+\partial_{\theta}\log p(\bm{y},\bm{\varphi}^{(q)};\bm{\theta}^{(q)})\partial_{\theta}\log p(\bm{y},\bm{\varphi}^{(q)};\bm{\theta}^{(q)})^{\prime}-G_{q-1}\right]$
$H_{q}$ $=$ $G_{q}-D_{q}D_{q}^{\prime}.$
At convergence, $-H_{q}^{-1}$ can be used to approximate the covariance matrix of the parameter estimates (Zhu and Lee,, 2002; Cai,, 2010), which are useful to derivate procedures for hypothesis testing for the different parameters of the model.
3 Simulation studies
To evaluate the behavior of the proposed estimation method, and to compare it with existing alternatives, we conducted several simulation studies. It is worth noticing that the GHQ approach does not allow to deal with a different number of observations per individual, which is possible with our SAEM-based approach and the gamlss package. Therefore, we present simulations with balanced data first. Additional simulations on unbalanced data are also provided in Appendix A, Supplementary Materials, in which the performance of SAEM on the unbalanced datasets is compared with the use of the GHQ algorithm on balanced datasets obtained from imputation, and with gamlss without imputation. Covariates significance analysis based on the LRT and the Wald test are also presented in Appendix B, Supplementary Materials.
3.1 Setup
We use two different settings for generating synthetic data under the ZIBR model (Equations 1 and 2). The parameters for each configuration were chosen as follows:
  • Setting 1: $a=b=-0.5$, $\alpha=\beta=0.5$, $\sigma_{1}=3.2$, $\sigma_{2}=2.6$, $\phi=6.4$.
  • Setting 2: $a=b=-0.5$, $\alpha=\beta=0.5$, $\sigma_{1}=0.7$, $\sigma_{2}=0.5$, $\phi=6.4$.
In the balanced scenario, for both Settings 1 and 2 the number of individuals $N=100$ will remain fixed, but the number of observations per individual $T_{i}$ will change, making $T_{i}=T$ with $T=3,5,10$. In addition, a variable $X$ is defined that mimics the concept of treatment and control groups, making $X=0$ for the first half of individuals and $X=1$ for the other half. Furthermore, we consider the same variable as covariate in both parts of the models, making $Z=X$.
For both Settings 1 and 2, $R=1000$ datasets were generated, and the SAEM estimation was implemented with $m=5$ chains and $(K_{1},K_{2})=(750,250)$, having therefore 1000 total iterations, with a starting point $\theta_{0}=(\phi_{0},a_{0},b_{0},\alpha_{0},\beta_{0},\sigma_{1,0},\sigma_{2,0})=(8,-0.3,-0.2,0.7,0.8,0.38,0.31)$.
3.2 Results
Figure 1: Estimated density of the parameters obtained by the SAEM algorithm, the GHQ method and the GAMLSS procedure on artificial balanced datasets simulated under Setting 1. The dotted vertical line represents the true value of the parameter.
Figure 2: Estimated density of the parameters obtained by the SAEM algorithm, the GHQ method and the GAMLSS procedure on artificial balanced datasets simulated under Setting 2. The dotted vertical line represents the true value of the parameter.
Table 1 shows the performance analysis of the two estimation methods for Settings 1 and 2 on balanced datasets, evaluated by bias $\left(\frac{1}{R}\sum_{r=1}^{R}\hat{\theta}^{r}-\theta\right)$, mean absolute error (MAE = $\left.\frac{1}{R}\sum_{r=1}^{R}|\hat{\theta}^{r}-\theta|\right)$ and root mean square error (RMSE = $\left.\sqrt{\frac{1}{R}\sum_{r=1}^{R}(\hat{\theta}^{r}-\theta)^{2}}\right)$.
Table 1: Summary statistics of the results obtained by SAEM algorithm, the GHQ procedure and the GAMLSS method on balanced data sets over 1000 simulation runs. For each parameter value and number of observations per individual, $T_{i}$, bold numbers indicate the lowest (absolute) value for each of bias, RMSE and MAE.
Parameter Value Bias RMSE MAE Bias RMSE MAE Bias RMSE MAE
SAEM GHQ GAMLSS
$a$0.5 $T_{i}=3$ 0.0064 0.5896 0.4629 -0.0010 0.6509 0.4734 0.1460 0.4348 0.3449
$T_{i}=5$ 0.0081 0.5626 0.4526 0.0044 0.5634 0.4514 0.1302 0.4281 0.3436
$T_{i}=10$ 0.0154 0.5369 0.4306 0.0221 0.5947 0.4749 0.1139 0.4269 0.3469
$\alpha$ 0.5 $T_{i}=3$ 0.0091 0.8277 0.6519 0.0215 0.9634 0.6587 -0.1343 0.5846 0.4626
$T_{i}=5$ 0.0046 0.8001 0.6340 0.0138 0.7926 0.6186 -0.1203 0.5822 0.4590
$T_{i}=10$0.0046 0.7533 0.6007 -0.0253 0.8363 0.6625 -0.1026 0.5827 0.4673
$b$0.5 $T_{i}=3$0.0414 0.4936 0.3970 0.0970 0.5570 0.4607 -0.0723 0.4744 0.3767
$T_{i}=5$0.0532 0.4511 0.3560 0.1752 0.5057 0.4317 -0.0829 0.4416 0.3478
$T_{i}=10$0.0552 0.4392 0.3512 0.3197 0.5155 0.4582 -0.0993 0.4375 0.3481
$\beta$ 0.5 $T_{i}=3$0.0384 0.6827 0.5372 -0.1442 0.7253 0.5788 -0.0493 0.6529 0.5122
$T_{i}=5$0.0216 0.6530 0.5092 -0.2532 0.6784 0.5544 -0.0526 0.6160 0.4809
$T_{i}=10$0.0275 0.6083 0.4825 -0.3828 0.5972 0.5202 -0.0660 0.5980 0.4704
$\sigma_{1}$ 3.2 $T_{i}=3$ 0.0284 0.6423 0.4961 0.0693 1.1306 0.5277 -0.7702 0.8400 0.7754
$T_{i}=5$ 0.0259 0.4920 0.3904 0.0631 0.5534 0.4238 -0.6914 0.7461 0.6932
$T_{i}=10$0.0012 0.3951 0.3170 0.0870 0.4433 0.3550 -0.5939 0.6428 0.5956
$\sigma_{2}$ 2.6 $T_{i}=3$0.1720 0.2979 0.2429 -0.3206 0.6612 0.4347 -0.0142 0.2478 0.1957
$T_{i}=5$0.1639 0.2779 0.2270 -0.4970 0.8065 0.5605 -0.0899 0.2425 0.1932
$T_{i}=10$0.1699 0.2635 0.2162 -0.8565 1.0651 0.8674 -0.1457 0.2480 0.2018
$\phi$ 6.4 $T_{i}=3$ 0.1301 1.2251 0.9465 -0.3696 2.0104 1.4482 4.6834 5.1779 4.6834
$T_{i}=5$ 0.0895 0.8059 0.6283 -0.9741 2.1842 1.4390 2.3698 2.6248 2.3713
$T_{i}=10$ 0.0645 0.4940 0.3931 -1.9106 2.8099 2.0202 1.1085 1.2564 1.1139
SAEM GHQ GAMLSS
$a$0.5 $T_{i}=3$ 0.0106 0.2101 0.1659 0.0026 0.2060 0.1623 0.0376 0.1935 0.1540
$T_{i}=5$ 0.0017 0.1813 0.1432 -0.0003 0.1780 0.1404 0.0317 0.1688 0.1340
$T_{i}=10$ 0.0033 0.1414 0.1144 0.0023 0.1380 0.1114 0.0276 0.1337 0.1080
$\alpha$ 0.5 $T_{i}=3$0.0120 0.2912 0.2283 -0.0040 0.2859 0.2243 -0.0385 0.2663 0.2102
$T_{i}=5$0.0060 0.2551 0.2025 -0.0032 0.2482 0.1973 -0.0349 0.2346 0.1870
$T_{i}=10$0.0042 0.1997 0.1607 -0.0029 0.1910 0.1528 -0.0282 0.1833 0.1468
$b$0.5 $T_{i}=3$ 0.0006 0.1348 0.1063 -0.0025 0.1331 0.1043 -0.0212 0.1409 0.1102
$T_{i}=5$0.0031 0.1099 0.0873 -0.0028 0.1096 0.0874 -0.0094 0.1116 0.0885
$T_{i}=10$0.0032 0.0922 0.0732 -0.0031 0.0922 0.0729 -0.0040 0.0928 0.0737
$\beta$ 0.5 $T_{i}=3$0.0089 0.1792 0.1430 -0.0049 0.1753 0.1403 0.0122 0.1829 0.1463
$T_{i}=5$0.0006 0.1536 0.1223 -0.0011 0.1516 0.1206 0.0050 0.1542 0.1226
$T_{i}=10$ 0.0015 0.1326 0.1060 0.0012 0.1308 0.1044 0.0012 0.1314 0.1052
$\sigma_{1}$ 0.7 $T_{i}=3$0.1448 0.3944 0.3239 -0.0535 0.3176 0.2501 -0.0845 0.2768 0.2164
$T_{i}=5$0.0394 0.2107 0.1605 -0.0275 0.1951 0.1499 -0.0643 0.1808 0.1373
$T_{i}=10$0.0192 0.1137 0.0910 -0.0169 0.1108 0.0883 -0.0489 0.1106 0.0877
$\sigma_{2}$ 0.5 $T_{i}=3$0.0812 0.1918 0.1407 -0.0542 0.1529 0.1139 0.2324 0.2481 0.2328
$T_{i}=5$0.0359 0.0991 0.0771 -0.0337 0.0963 0.0749 0.1453 0.1610 0.1460
$T_{i}=10$0.0193 0.0629 0.0499 -0.0190 0.0624 0.0496 0.0715 0.0891 0.0754
$\phi$ 6.4 $T_{i}=3$0.0920 1.2570 1.0054 0.0031 1.1890 0.9402 7.4381 8.0589 7.4381
$T_{i}=5$0.0700 0.7283 0.5820 -0.0620 0.7214 0.5767 3.3718 3.6059 3.3718
$T_{i}=10$0.0463 0.4423 0.3518 -0.0472 0.4409 0.3510 1.3984 1.5107 1.4002
Analyzing these results globally, and the global estimates distributions in Figures 1 and 2, we can see how SAEM estimates are always centered (except for a small bias for $\sigma_{2}$ in Setting 1), whereas GHQ and GAMLSS methods present strong biases or bimodality for different parameters: $b$, $\beta$, $\sigma_{2}$ and $\phi$ for GHQ in Setting 1 and $\sigma_{1}$ (only Setting 1), $\sigma_{2}$ and $\phi$ (both settings) for GAMLSS. A thorough examination of these results reveals that the SAEM estimation achieves optimal bias performance in Setting 1. In Setting 2, the GHQ method exhibits slightly lower bias values. It merits attention, however, that GAMLSS attains the lowest root mean square error (RMSE) and mean absolute error (MAE) values in many parameters across both settings. It is worth noticing, nevertheless, that the GAMLSS estimates for $\phi$ are extremely biased in both settings. These two situations are common in estimators obtained by quasi-likelihood methods, as indicated by Nelder and Lee, (1992).
A remarkable property of SAEM is its tendency to exhibit a consistent decrease in the error measures as the number of observations per individual increases, a phenomenon not consistently observed in the other analyzed alternatives. In Setting 1, where the values of the variance parameters are higher, it is evident that the worst GHQ results are obtained in the estimation of the parameters associated with the Beta part of the model ($b,\beta,\sigma_{2}$ and $\phi$). This component has a complex functional form that could be incorrectly approximated when integrating by numerical methods. The SAEM approach is advantageous in this regard. In the context of GAMLSS, the estimation of the parameter controlling the overdispersion of the data, i.e., $\phi$, is particularly challenging (as well as the estimation of $\sigma_{2}$ in Setting 2). According to Stasinopoulos et al., (2017), by defining ZIBR as a model with several random effects, GAMLSS estimates the variance components with a method prone to generating biases. Given that overdispersion is a common feature in real microbiota data, the GAMLSS procedure may not be suitable for modeling this phenomenon.
A more detailed analysis of the Figures shows that for Setting 1 the estimated densities of $a$ and $\alpha$ are very similar in all methods, while the other parameters show marked differences. In the case of GHQ, the results for $b$ and $\beta$ are very skewed, while the distribution of $\sigma_{2}$ and $\phi$ shows bimodal behavior. This is due to the poor approximation of the log-likelihood, which leads to an erroneous estimation of these parameters. A review of the GAMLSS results reveals that the distribution of $\sigma_{1}$ is significantly deviated from the true parameter, while the distribution of $\phi$ exhibits considerable variability when observations per individual are limited. Figure 2, on the other hand, shows that for Setting 2 SAEM and GHQ are practically equivalent in the estimates densities, while GAMLSS differs only in that it performs poorly for $\sigma_{2}$ and $\phi$. This indicates that the minor errors differences between the methods in favor of GHQ for this simulation scenario (Table 1) are not significant.
Regarding the execution times of the routines, GHQ and GAMLSS are faster than SAEM in the 3 cases considered in both Settings. On a computer with Intel Core i7-13700HX processor at 2.10 GHz, for $T_{i}=3\ (5,10)$, on average, GAMLSS takes $0.94\ (1.03,1.20)$ seconds, GHQ takes $1.77\ (3.39,8.54)$ seconds, while SAEM takes $9.86\ (14.14,25.30)$ seconds.
4 Case studies
In this section we demonstrate the use of the proposed inference framework on a publicly available microbiome study and we will focus on the capabilities of the SAEM algorithm to detect changes in the presence of bacterial taxa in response to treatments. The study of another set of microbiome data can be found in Appendix C of the Supplementary Materials. As we mentioned in the previous section, given that the ZIBR model has more than one random effect, GAMLSS uses an estimation method that can cause a bias in the estimation of the variance parameters, and also does not allow the evaluation of the tests based on the calculation of the log-likelihood that we will use in the following section. For this reason, the GAMLSS approach will not be applied to the real data considered below.
4.1 Inflammatory bowel disorder pediatric study
The data used in this section come from a study to verify the effectiveness of treatments in pediatric inflammatory bowel disorder (IBD) patients (Lewis et al.,, 2015). This study includes information from 90 children subjected to three different types of therapy: anti-TNF treatment (TNF: tumor necrosis factor), exclusive enteral nutrition (EEN) and partial enteral nutrition with an ad lib diet (PEN). After filtering the data to discard low sequencing depth samples, low abundant genus and taxa with a proportion of zeros higher than 0.9 or lower than 0.1, the information from 18 bacterial genera measured at 4 different time points for each one of the 59 individuals (47 anti-TNF and 12 EEN) remained for the analysis. Figure 3 shows the average composition of the intestinal microbiome of the subjects in both groups and its evolution over time.
Figure 3: Average gut microbiome composition of the treatment groups (anti-TNF and EEN) over observation week
The objective of the study is to verify if the different treatments influence the presence of the different bacterial taxa in the samples, controlling for time and initial abundance. In addition, we want to compare if the results obtained by the SAEM algorithm differ from those obtained through the GHQ procedure implemented in the ZIBR package. The initial values for SAEM were the estimates found by the GHQ method, for each model corresponding to each bacterial taxon. Given this choice of initial values, we used $m=5$ Markov chains and $(K_{1},K_{2})=(375,125)$ iterations, as well as 500 simulated values for the log-likelihood calculation by Importance Sampling. The p-values obtained through the LRT were adjusted using the Benjamini-Hochberg process (Benjamini and Hochberg,, 1995) to decrease the false discovery rate (FDR) and the full values are presented in Table 8, Appendix D of Supplementary Materials.
Figure 4: Bacterial taxa in which the treatment (anti-TNF vs. EEN) have a statistical effect in abundance identified by SAEM and GHQ
After model fitting, the GHQ method detected 11 bacterial taxa in which the treatment influenced the abundances, while SAEM managed to identify 12 taxa, all those identified by GHQ in addition to Escherichia with FDR=5%, as shown in Figure 4.
A more detailed analysis of the Escherichia data shows that the influence of treatment is greater on the frequency of presence of this bacterium in individuals than on the level of abundance once its presence is confirmed (Figure 5 in Appendix D). This is confirmed by the LRT results for the significance of the treatment in the calculation of the probability $p_{it}$ (FDR p-value 0.03) compared to those of the Beta component of the abundance $u_{it}$ (FDR p-value 0.80), and by the Wald test (Table 9, Appendix D). These results show that at the 5% significance level the treatment is significant in the logistic part but not in the Beta part, proving that the definition of the ZIBR model and the combination with the SAEM estimation allows the increase of the ability to detect the influence of a given treatment defined. A detailed figure with the convergence behavior of the estimators across iterations is shown in Figure 6, Appendix D of the supplementary materials.
The role of Escherichia in IBD is well documented. There is evidence (Baldelli et al.,, 2021) that the accumulation of Escherichia coli and other strains of Escherichia in the intestine is related to inflammatory processes, and other works (Mirsepasi-Lauridsen et al.,, 2019) suggest that a combination of antibiotic and dietary treatments is capable of controlling overproliferation of E. coli in the digestive system and also reducing the symptoms of IBD, allowing to infer a correlation between these two events.
5 Conclusions and discussion
In this article we have developed an exact maximum likelihood estimation strategy for the ZIBR model for the analysis of longitudinal compositional microbiome data using the SAEM algorithm. We have also proposed a method for calculating the log-likelihood of the model which allows to obtain information criteria for the model, and approximations of the estimators standard errors, which is not possible under the alternative estimation method based on Gauss-Hermite quadrature likelihood approximation. Moreover, despite the capacity of the GAMLSS approach to estimate the ZIBR model parameters and calculate the standard errors of these estimates, the SAEM method exhibits superior performance controlling the type I error of the Wald test for the significance of parameters. Although the results obtained by SAEM conform to the expected theoretical properties of standard errors, it must be noted that the method can still be improved, since, as it depends on a stochastic approximation, convergence towards coherent values for the Fisher information matrix is not guaranteed, which could introduce bias in the estimation of the standard errors. We are confident that these details can be enhanced in further developments.
Another advantage of the proposed estimation method is its ability on handling unbalanced data, a scenario that can occur both due to the design of the experiment itself and due to external factors, such as to individuals dropping out during the study. This aspect was not considered in the development of the original estimation method for ZIBR, so comparisons of performance with our method cannot be established in this scenario, unless data interpolation is performed before using the GHQ method. The GAMLSS method, another option that allows for handling unbalanced data, has been found to show errors in the estimation of certain important parameters of the model, even when its performance in the rest is adequate. It should be emphasized that unbalanced data is a fairly common situation in medical experiments, in which multiple factors influence patients abandoning the follow-up. This could be one of the reasons contributing to the high non-publication rate in many medical studies, which according to certain sources could be close to 50% (Chan et al.,, 2014). Therefore, developing analysis methods that can deal accurately with unbalanced data is of great interest.
The definition of the ZIBR model used throughout this work corresponds to the one originally proposed by Chen and Li, (2016) and implemented in the ZIBR package for R statistical software. However, there are possibilities for modification of this definition that have been discussed. One of them is the use of random effects for more than one covariate, an aspect that has been already incorporated in the implementation used in this article. Another possibility is the inclusion of cross correlations in the random effects, proposing a different structure in the variance of these effects. Liu et al., (2019) mention that this inclusion could alter the results for tests on covariates, detecting significance where a simpler structure would not detect it. Although this approach has not been implemented here, the SAEM algorithm could be easily modified to serve this purpose.
In the field of human microbiome analysis, other models have been proposed in addition to ZIBR. Among the most important are ZIBR-SRE (Han et al.,, 2021), an extension of ZIBR which considers the compositional nature of microbiota data; zero-inflated Gaussian mixed models (ZIGMM) (Zhang et al.,, 2020), which in addition to managing the overabundance of zeros can work with both proportion data and counts; and the negative binomial mixed model (NBMM) (Zhang et al.,, 2018), which allows the specification of more general variance structures and also be modified to deal with zero inflation. It seems interesting to implement the SAEM algorithm to these models and study its potential benefits in estimation.
Finally, an interesting extension of this work would be to obtain the Restricted Maximum Likelihood (REML) estimates, a known method for reducing bias in the estimation of variance components in mixed effects models (Meza et al.,, 2007), using the Harville’s approach, i.e., integrating out the fixed effects, via the SAEM algorithm. We expect that, in the context of longitudinal models on microbiome data, this could improve the results obtained through ML estimation.
Supplementary materials
The supplementary materials include extended simulation studies, tables and figures referenced in Sections 3 and 4 and an additional case study. R codes to implement the routines of the SAEM estimation and to reproduce the statistical analysis of Section 4 are available at https://github.com/jbarrera232/saem-zibr.
Declaration of conflicting interests
The authors declared no potential conflicts of interest with respect to the research, authorship and/or publication of this article.
Funding
The work of the first and second author was supported by ANID MATH-AmSud Project AMSUD 230032-SMILE. The work of the first author was also supported by ANID Becas/Doctorado Nacional 21231659. The third author gratefully acknowledges the support of grants PID2021-123592OB-I00 funded by MCIN/AEI/10.13039 /501100011033 and by ERDF - A way of making Europe, and TED2021-129316B-I00 funded by MCIN/AEI/10.13039 /501100011033 and by the European Union NextGenerationEU/PRTR.